One side of the rectangle is 7cm shorter than five times the other side. Find the length of the sides if the perimeter of the rectangle is 118 cm.
Answer by ReadingBoosters(3246)

2(x)+ 2(5x-7) = 118
2x + 10x - 14 = 118
12x = 132
x = 11 cm
5(11)-7 = 48 cm
sides are 11 cm and 48 cm
Proof
2(11) + 2(48) = 118
..................
Delighted to help.
-Reading Boosters
Wanting for others what we want for ourselves.